# Net Charge Accretion in Magnetized Kerr Black Holes

### v2

@@ I Introduction
When a rotating black hole is immersed in an environmental magnetic field, the strong gravitational field distorts the electromagnetic field, creating an electric field – which may be strong in the close vicinity of the BH. Such an induced electric field may potentially lead to various astrophysical effects (Ruffini2023; Levin2018; King2021).
About half a century ago, Wald investigated the case of a Kerr BH immersed in an asymptotically uniform magnetic field aligned with the BH’s axis of rotation (where the EM field is assumed to be too small to affect the background metric) (Wald1974). He found that in this case, a (current-free) four potential $A_{\mu}\mathopen{}\mathclose{{\left(x}}\right)$ can be directly constructed from the two Killing fields of the Kerr background. This construction leads to a two-parameter family of solutions to Maxwell’s equations. One of them is the asymptotic magnetic field magnitude $B_{0}$, and the second parameter, $Q$, is the BH’s charge.
@@
In his paper (Wald1974), Wald also investigated the expected value of the saturation charge under the assumption that the so-called injection energy of charged particles into the BH should vanish at $Q=Q_{\mathrm{sat}}$. This assumption led to a simple, universal value for the saturation charge $Q_{\mathrm{sat}}=2B_{0}J\equiv Q_{\mathrm{w}}$ where $J$ denotes the BH’s angular momentum. As a technical remark, note that $Q_{\mathrm{w}}$ is positive if the angular momentum and the asymptotic magnetic field are co-directed, and negative otherwise. Throughout this paper, without loss of generality, we shall assume that $Q_{\mathrm{w}}>0$. We shall often refer to negatively charged particles as attracted and to positively charged ones as repelled.
@@
Motivated by the desire to better understand the saturation charge $Q_{\mathrm{sat}}$, we set out to develop an accretion model as natural and simple as possible, which will allow us to quantitatively explore the absorption rates in terms of the corresponding cross sections – and thereby, to determine whether there actually is charge accretion balance or imbalance at $Q=Q_{\mathrm{w}}$. In our setup, there are two twin uniform fluxes of unbound positively and negatively charged particles coming from infinity along the magnetic field lines. These two fluxes are fully equivalent, in the sense that they share the same flux density, and the positive and negative charges have the same mass $m$, the same charge magnitude $\mathopen{}\mathclose{{\left|e}}\right|$, and the same asymptotic incoming velocity – they only differ by the sign of the particles’ charge (a more detailed description of our accretion model is given in the next section). We assume that the fluxes are very dilute, such that individual charges do not affect each other’s motion. In this situation, assuming equal fluxes, the issue of charge accretion balance (or imbalance) boils down to the question of whether the two absorption cross sections are equal or not – which is the question on which we focus throughout this paper.
@@
Naively, one might expect to find a single continuous domain of absorption that extends from $b=0$ up to a certain critical value $b_{\mathrm{crit}}$ beyond which no orbit would be absorbed (in which case the cross section would be $\pi\mathopen{}\mathclose{{\left(b_{\mathrm{crit}}}}\right)^{2}$). Indeed, there is a central, continuous, absorption domain from $b=0$ up to a certain value $b_{1}$; but to our surprise, we found that at $b>b_{1}$, the family of orbits has a fractal structure. In particular, in the range $b>b_{1}$, there appear to exist an infinite set of individual absorption (as well as non-absorption) domains. We note that such a fractal behavior has been observed in magnetized Schwarzschild (Frolov2013) and Kerr (Levin2018; Adari2023) BHs. nevertheless, this fractal behavior has not been observed or analyzed so far in the context of a family of unbound orbits coming from infinity.
@@
In order to investigate charge accretion imbalance, we proceed in two stages. We start by choosing a specific point in the space of our model’s parameters – namely specific values of $\frac{J}{M^{2}}$, $E$ and $B_{0}$ – for which the analysis becomes especially simple, and demonstrate imbalance in this specific case. Then, in the next stage, we extend the analysis to the entire range of $B_{0}$ and show that charge accretion imbalance is guaranteed to occur beyond a certain value of $B_{0}$. Furthermore, this imbalance diverges at the limit of large $B_{0}$.
In Sec. II we review Wald’s solution for $A_{\mu}\mathopen{}\mathclose{{\left(x}}\right)$ and use it in our accretion analysis. As already mentioned above, in our model a charged particle comes in from infinity along the magnetic field lines with a certain energy $E>1$ and impact parameter $b$. Due to the scale invariance of GR, the relevant model’s parameters reduce to the BH’s specific spin $\alpha\equiv\frac{J}{M^{2}}$, the magnetic field strength parameter $\varepsilon\equiv\frac{e}{m}B_{0}M$, as well as $E$ and $b$. We write down the particles’ EOM, and relate the impact parameter $b$ to the conserved angular momentum $L$.
@@ II.1 Wald solution
It has been shown by Wald (Wald1974) that the vector potential created by a stationary and axisymmetric Kerr BH, placed in an asymptotically uniform magnetic field which is asymptotically aligned with the axis of rotation, can be constructed using the timelike and axial Killing vectors $\eta^{\mu}\equiv g_{t}^{\mu}$ and $\psi^{\mu}\equiv g_{\phi}^{\mu}$ respectively, where $g_{\mu\nu}$ is the spacetime metric. Wald’s potential is given by
@@
where $M$ and $J$ are respectively the mass and angular momentum of the BH, and $B_{0}$ is the magnetic field magnitude at infinity. The parameter $Q$ is the charge of the BH, assumed to be too small to have an effect on the metric [^1] The effect of the electromagnetic field (associated with the BH’s charge $Q$) on the metric is expected to be negligible for known astrophysical examples (See (Wald1974)).
@@
In this paper, unless explicitly stated otherwise, we’ll choose our units such that the mass of the BH, $M$, is equal to $1$ (in addition to $G=c=1$). This allows us to rewrite the four-potential as
@@ II.2 Equations of motion
Throughout this paper we assume that the particle is moving in the Kerr metric
@@ II.3 Defining the physical system
All particles are assumed to be of the same mass $m$ and the same charge magnitude which is either +e+e or e-e. The (asymptotically uniform) influx of charges of both signs is assumed to be equal, and all particles have the same energy $E$. We are primarily interested in the absorption cross sections $\sigma_{\pm}$ for charges of both signs. These two cross sections depend on the trajectory parameter $E$, as well as the background parameters $\alpha$ and $\varepsilon$. The overall charge accretion will be balanced if and only if $\sigma_{+}=\sigma_{-}$.
@@ III.1 Trajectory types
For the purpose of investigating the cross section for absorption, it is useful to categorize particle trajectories according to their ultimate end point. This results in the definition of three different trajectory types: 1) Falling trajectories, which eventually cross the event horizon. For example, since the electric field along the symmetry axis vanishes ($F_{tr}=0$), an incoming charged particle along the axis of symmetry will necessarily fall into the BH. 2) Escaping trajectories, that eventually go to infinity. For example, a particle with a large enough impact parameter will inevitably escape to infinity, only weakly interacting with the BH and the electric potential. 3) Trapped trajectories, that neither fall into the BH or escape to infinity. For example, a particle moving in an equatorial circular orbit. The existence and categorization of such trajectories have been explored in previous studies (Liu2018; Sun2021; Zaharani2014; Kopacek2018; Aliev2002; Frolov2010). In Subsec. III.2 we discuss another type of trapped trajectory which is of importance to the analysis of domains of accretion.
The fact that the $b=0$ orbit is always absorbed, combined with the basic continuity of the solutions to the EOM with respect to the initial conditions, imply that in some neighborhood of $b=0$ all trajectories are falling. On the other hand, for large enough $b$ (with $\varepsilon$, $E$ and $\alpha$ remaining fixed), all trajectories are obviously escaping. We accordingly define $b_{1}$ as the maximal $b$ value up to which all trajectories are falling, and $b_{2}$ as the minimal $b$ value beyond which all trajectories are escaping. In the range $b_{1}<b<b_{2}$ trajectories can either fall, escape, or be trapped. Due to its nature, further explained in Sec. VI, we refer to this intermediate range as the fractal domain.
@@ III.3 Universal limit of large $\varepsilon$
It is convenient to investigate this limit at the level of the EOM by reparametrizing the orbit with $\tau^{\prime}=\varepsilon\tau$ instead of proper time $\tau$, and correspondingly, defining the rescaled four-velocity $u^{\prime\mu}=\frac{dx^{\mu}}{d\tau^{\prime}}$. Importantly, substituting these new variables in Eq. (4) yields well-behaved and effective EOM, as will be shown below. We refer to these EOM as “universal” as they are independent of both $\varepsilon$ and $E$, as we’ll now discuss.
@@
$u^{\prime\mu}u^{\prime}_{\mu}=0$ (20)
where $\tilde{F}\equiv F/\varepsilon$ is the rescaled electromagnetic tensor, which depends only on the spin parameter $\alpha$. Thus, the EOM in the universal limit portrays a picture of a massless charged particle with zero energy and fixed angular momentum, moving in a magnetized Kerr background with finite field strength $\tilde{F}$.
@@
$b_{1}^{\mathrm{univ}}\equiv\lim_{\varepsilon\to-\infty}b_{1}\mathopen{}\mathclose{{\left(\varepsilon}}\right)$
can be found numerically using the same method presented in Subsec. III.2, and it presents the same qualitative behavior as described therein. Fig. 1 shows numerical results for the universal critical impact parameter $b_{1}^{\mathrm{univ}}$ as a function of spin $\alpha$, computed using the universal EOM (19). As expected, for any fixed $\alpha$, $b_{1}\mathopen{}\mathclose{{\left(\varepsilon}}\right)$ converges to the corresponding critical impact parameter $b_{1}^{\mathrm{univ}}$ as $\varepsilon\rightarrow$-$\infty$, and this limiting parameter is $E$-independent. This is illustrated in Fig. 2.
@@ IV.1 Energetically allowed and forbidden regions
Note that the function $n$ depends only on the coordinates $r$ and $\theta$ (in addition to its dependence on the system parameters $\alpha$ and $\varepsilon$, and the orbit parameters $E$ and $b$). The particles’ orbits must satisfy
@@
On the axis of symmetry, namely at $\theta=0$, the denominator of $n$ in Eq. (23) vanishes. If $n_{\mathrm{u}}>0$ on the axis, then $n$ diverges to $\infty$ there, meaning that the axis and its neighborhood are forbidden in such a case. Therefore, in order for the axis and its neighborhood to be energetically allowed, Eq. (26) must hold at $\theta=0$ as well.
@@ IV.2 Globally connecting and disconnecting points
On the other hand, all orbits with sufficiently large $b$ values are forbidden for absorption. To see this, note that (i) $n_{\mathrm{u}}$ is a polynomial of second order in $b^{2}$ whose coefficients are themselves polynomial in $r$ and $\sin^{2}\theta$; (ii) the coefficient of the highest order in $b$ (namely, $b^{4}$) is $\varepsilon^{2}r\mathopen{}\mathclose{{\left(r-2}}\right)+\varepsilon^{2}\alpha^{2}\mathopen{}\mathclose{{\left(1-\sin^{2}\theta}}\right)$, which is always positive for $r>2$. Choose now any “circle” of $r=const>2$ and evaluate the sign of $n_{\mathrm{u}}$ on this curve. Both the coefficients of the $b^{2}$ and $b^{0}$ terms in $n_{\mathrm{u}}$, being polynomials in $\sin^{2}\theta$, are bounded from above at that $r$ value. Since the coefficient of highest order term $b^{4}$ is positive, it is obvious that at sufficiently large $b$, on this constant-$r$ curve $n_{\mathrm{u}}>0$, meaning that this circle is entirely in a forbidden region. In other words, there is no allowed region connecting the horizon to infinity.
From the above, it follows that there must exist a value of $b$ that marks a transition (with increasing $b$) between an allowed $b$-range and a forbidden $b$-range. We call such a transition point (on the $b$ axis) a globally disconnecting point. In principle, there may also exist a point of transition in the other direction (namely, from a forbidden range to an allowed one) which we call a globally connecting point.

### v3

@@ I Introduction
When a rotating black hole (BH) is immersed in an environmental magnetic field, the strong curvature distorts the electromagnetic field, creating an electric field – which may be strong in the close vicinity of the BH. Such an induced electric field may potentially lead to various astrophysical effects (Ruffini2023; Levin2018; King2021).
About half a century ago, Wald investigated the case of a Kerr BH immersed in an asymptotically uniform magnetic field aligned with the BH’s axis of rotation (where the electromagnetic field is assumed to be too small to affect the background metric) (Wald1974). He found that in this case, a (current-free) four potential $A_{\mu}\mathopen{}\mathclose{{\left(x}}\right)$ can be directly constructed from the two Killing fields of the Kerr background. This construction leads to a two-parameter family of solutions to Maxwell’s equations. The first parameter is the asymptotic magnetic field magnitude $B_{0}$, and the second parameter, $Q$, is the BH’s charge.
@@
In his paper (Wald1974), Wald also investigated the expected value of the saturation charge under the assumption that the so-called injection energy of charged particles into the BH should vanish at $Q=Q_{\mathrm{sat}}$. This assumption led to a simple, universal value for the saturation charge $Q_{\mathrm{sat}}=2B_{0}J\equiv Q_{\mathrm{w}}$ where $J$ denotes the BH’s angular momentum. As a technical remark, note that $Q_{\mathrm{w}}$ is positive if the BH’s angular momentum and the asymptotic magnetic field are co-directed, and negative otherwise. Throughout this paper, without loss of generality, we shall assume that $Q_{\mathrm{w}}>0$. Correspondingly, we shall often refer to negatively charged particles as attracted and to positively charged ones as repelled, which will be denoted with superscripts “-” and “+” respectively.
@@
The problem of charge accretion into a magnetized Kerr BH has previously been considered in Rueda2024 (under the simplifying assumption that charged particles move along the magnetic field lines, even in the BH’s vicinity). Their results seem to differ from Wald’s picture in several respects. In particular, the authors contemplate the possibility that the BH charge oscillates rather than approaches a steady state. Another question mark regarding Wald’s picture was raised in Li2000 where it was suggested that the system’s total electromagnetic-field energy may be a more relevant criterion than the injection energy of individual charges. The charge that minimizes the total electromagnetic energy is found there to differ from Wald’s charge. In addition, although the system considered in Adari2023 differs from the one analyzed by Wald (they consider a boosted magnetized BH), we find it relevant to mention the authors’ viewpoint that detailed particle dynamics (rather than mere energetic arguments) should determine the BH’s charge.
Motivated by the desire to better understand the saturation charge $Q_{\mathrm{sat}}$, in this work we set out to develop an accretion model as natural and simple as possible, which will allow us to quantitatively explore the absorption rates in terms of the corresponding cross sections – and thereby, to determine whether there actually is charge accretion balance or imbalance at $Q=Q_{\mathrm{w}}$. In our setup, there are two twin uniform fluxes of unbound positively and negatively charged particles coming from infinity along the magnetic field lines. These two fluxes are fully equivalent, in the sense that they share the same flux density, and the positive and negative charges have the same mass $m$, the same charge magnitude $\mathopen{}\mathclose{{\left|e}}\right|$, and the same asymptotic incoming velocity – they only differ by the sign of the particles’ charge (a more detailed description of our accretion model is given in the next section). We assume that the fluxes are very dilute, such that individual charges do not affect each other’s motion. In this situation, assuming equal fluxes, the issue of charge accretion balance (or imbalance) boils down to the question of whether the two absorption cross sections are equal or not – which is the question on which we focus throughout this paper.
@@
Naively, one might expect to find a single continuous domain of absorption that extends from $b=0$ up to a certain critical value $\tilde{b}$ beyond which no orbit would be absorbed (in which case the cross section would be $\pi\tilde{b}^{2}$). Indeed, there is a central, continuous, absorption domain from $b=0$ up to a certain value $b_{1}$; but to our surprise, we found that at $b>b_{1}$, the family of orbits has a fractal structure. In particular, in the range $b>b_{1}$, there appear to exist an infinite set of individual absorption (as well as non-absorption) domains. We note that such a fractal behavior has been observed in magnetized Schwarzschild (Frolov2013) and Kerr (Levin2018; Adari2023) BHs. nevertheless, this fractal behavior has not been observed or analyzed so far in the context of a family of unbound orbits coming from infinity.
@@
In order to investigate charge accretion imbalance, we proceed in two stages. We start by choosing a specific point in the space of our model’s parameters – namely specific values of $\frac{J}{M^{2}}$, $E$ and $B_{0}$ – for which the analysis becomes especially simple, and demonstrate imbalance in this specific case. Then, in the next stage, we extend the analysis to the entire range of $B_{0}$ and show that charge accretion imbalance is guaranteed to occur beyond a certain value of $B_{0}$. Moreover, the “imbalance ratio” $\frac{\sigma^{-}}{\sigma^{+}}$ diverges at the limit of large $B_{0}$ [^1] As a matter of fact, the magnitude of the magnetic field will later be quantified by means of the dimensionless parameter $\varepsilon$ defined below, rather than $B_{0}$ itself..
In Sec. II we review Wald’s solution for $A_{\mu}\mathopen{}\mathclose{{\left(x}}\right)$ and use it in our accretion analysis. As already mentioned above, in our model a charged particle comes in from infinity along the magnetic field lines with a certain energy $E>1$ and impact parameter $b$. Due to the scale invariance of general relativity, the relevant model’s parameters reduce to the BH’s specific spin $\alpha\equiv\frac{J}{M^{2}}$, the magnetic field strength parameter $\varepsilon\equiv\frac{e}{m}B_{0}M$, as well as $E$ and $b$. We write down the particles’ EOM, and relate the impact parameter $b$ to the conserved angular momentum $L$.
@@ II.1 Wald solution
It has been shown by Wald (Wald1974) that the source-free vector potential induced by a stationary and axisymmetric Kerr BH, placed in an asymptotically uniform magnetic field which is asymptotically aligned with the axis of rotation, can be constructed using the timelike and axial Killing vectors $\eta^{\mu}\equiv g_{t}^{\mu}$ and $\psi^{\mu}\equiv g_{\phi}^{\mu}$ respectively, where $g_{\mu\nu}$ is the spacetime metric, and $t$ and $\phi$ are respectively the standard temporal and azimuthal Boyer-Lindquist coordinates, see Eq. (5). Wald’s potential is given by
@@
where $M$ and $J$ are respectively the mass and angular momentum of the BH, and $B_{0}$ is the magnetic field magnitude at infinity. The parameter $Q$ is the charge of the BH, assumed to be too small to have an effect on the metric [^2] The effect of the electromagnetic field (associated with the BH’s charge $Q$) on the metric is expected to be negligible for known astrophysical examples (see (Wald1974))..
@@
In this paper, unless explicitly stated otherwise, we will choose our units such that the mass of the BH, $M$, is equal to $1$ (in addition to $G=c=1$). This allows us to rewrite the four-potential as
@@ II.2 Equations of motion
Throughout this paper we assume that the particle is moving in the Kerr metric, given in Boyer-Lindquist coordinates (with $M$ set to 1) by
@@ II.3 Defining the physical system
All particles are assumed to be of the same mass $m$ and the same charge magnitude which is either +e+e or e-e. The (asymptotically uniform) influx of charges of both signs is assumed to be equal, and all particles have the same energy $E$. We are primarily interested in the absorption cross sections $\sigma^{\pm}$ for charges of both signs. These two cross sections depend on the trajectory parameter $E$, as well as the background parameters $\alpha$ and $\varepsilon$. The overall charge accretion will be balanced if and only if $\sigma^{+}=\sigma^{-}$.
@@ III.1 Trajectory types
For the purpose of investigating the cross section for absorption, it is useful to categorize particle trajectories according to their ultimate end point. This results in the definition of three different trajectory types: 1) Falling trajectories, which eventually cross the event horizon. For example, since the electric field along the symmetry axis vanishes ($F_{tr}=0$), an incoming charged particle along the axis of symmetry will necessarily fall into the BH. 2) Escaping trajectories, that eventually go to infinity. For example, a particle with a large enough impact parameter will inevitably escape to infinity, only weakly interacting with the BH and the electric potential. 3) Trapped trajectories, that neither fall into the BH nor escape to infinity. For example, a particle moving in an equatorial circular orbit. The existence and categorization of such trajectories have been explored in previous studies (Liu2018; Sun2021; Zaharani2014; Kopacek2018; Aliev2002; Frolov2010). In Subsec. III.2 we discuss another type of trapped trajectory which is of importance to the analysis of domains of accretion.
The fact that the $b=0$ orbit is always absorbed, combined with the basic continuity of the solutions to the EOM with respect to the initial conditions, imply that in some neighborhood of $b=0$ all trajectories are falling. On the other hand, for large enough $b$ (with $\varepsilon$, $E$ and $\alpha$ remaining fixed), all trajectories are obviously escaping. We accordingly define $b_{1}$ as the maximal $b$ value up to which all trajectories are falling, and $b_{2}$ as the minimal $b$ value beyond which all trajectories are escaping. In the range $b_{1}<b<b_{2}$ trajectories can either fall, escape, or be trapped. We further discuss this complex behaviour (with specific regard to the fractal domain of falling orbits) in Sec. VI.
@@ III.3 Universal limit of large $\varepsilon$
It is convenient to investigate this limit at the level of the EOM by reparametrizing the orbit with $\tau^{\prime}=\varepsilon\tau$ instead of proper time $\tau$, and correspondingly, defining the rescaled four-velocity $u^{\prime\mu}=\frac{dx^{\mu}}{d\tau^{\prime}}$. Importantly, substituting these new variables in Eq. (4) yields well-behaved and effective EOM, as will be shown below. We refer to these EOM as “universal” as they are independent of both $\varepsilon$ and $E$, as we will now discuss.
@@
$u^{\prime\mu}u^{\prime}_{\mu}=0~,$ (20)
where $\tilde{F}\equiv F/\varepsilon$ is the rescaled electromagnetic tensor, which depends only on the spin parameter $\alpha$. Thus, the EOM in the universal limit portray a picture of a massless charged particle with zero energy and fixed angular momentum, moving in a magnetized Kerr background with finite field strength $\tilde{F}$.
@@
$b_{1}^{\mathrm{univ}}\mathopen{}\mathclose{{\left(\alpha}}\right)\equiv\lim_{\varepsilon\to-\infty}b_{1}\mathopen{}\mathclose{{\left(\varepsilon;\alpha,E}}\right)$
can be found numerically using the same method presented in Subsec. III.2, and it presents the same qualitative behavior as described therein. Fig. 1 shows numerical results [^5] Note that in Fig. 1, as well as in Figs. 2,6 and 7, the actual numerically calculated data are represented by points on the graphs while the curves connecting them are merely interpolations added for better visual illustration. for the universal critical impact parameter $b_{1}^{\mathrm{univ}}$ as a function of spin $\alpha$, computed using the universal EOM (19). As expected, for any fixed $\alpha$, $b_{1}\mathopen{}\mathclose{{\left(\varepsilon;\alpha,E}}\right)$ converges to the corresponding critical impact parameter $b_{1}^{\mathrm{univ}}$ as $\varepsilon\rightarrow$-$\infty$, and this limiting parameter is $E$-independent. This is illustrated in Fig. 2.
@@ IV.1 Energetically allowed and forbidden regions
Note that the function $n$ depends only on the coordinates $r$ and $\theta$ (in addition to its dependence on the system parameters $\alpha$ and $\varepsilon$, and the orbit parameters $E$ and $b$). As mentioned, the particles’ orbits must satisfy
@@
On the axis of symmetry, namely at $\theta=0$ (or equivalently $\theta=\pi$), the denominator of $n$ in Eq. (23) vanishes. If $n_{\mathrm{u}}>0$ on the axis, then $n$ diverges to $\infty$ there, meaning that the axis and its neighborhood are forbidden in such a case. Therefore, in order for the axis and its neighborhood to be energetically allowed, Eq. (26) must hold at $\theta=0$ as well.
@@ IV.2 Globally connecting and disconnecting points
On the other hand, all orbits with sufficiently large $b$ values are forbidden for absorption. To see this, note that (i) $n_{\mathrm{u}}$ is a polynomial of second order in $b^{2}$ whose coefficients are themselves polynomial in $r$ and $\sin^{2}\theta$; (ii) the coefficient of the highest order in $b$ (namely, $b^{4}$) is $\varepsilon^{2}r\mathopen{}\mathclose{{\left(r-2}}\right)+\varepsilon^{2}\alpha^{2}\mathopen{}\mathclose{{\left(1-\sin^{2}\theta}}\right)$, which is always positive for $r>2$. Choose now any “circle” of $r=const>2$ and evaluate the sign of $n_{\mathrm{u}}$ on this curve. Both the coefficients of the $b^{2}$ and $b^{0}$ terms in $n_{\mathrm{u}}$, being polynomials in $\sin^{2}\theta$, are bounded from above at that $r$ value. Since the coefficient of the highest order term $b^{4}$ is positive, it is obvious that at sufficiently large $b$, $n_{\mathrm{u}}>0$ is satisfied everywhere on this constant-$r$ curve, meaning that this circle is entirely in a forbidden region. In other words, there is no allowed region connecting the horizon to infinity.
From the above, it follows that there must exist a value of $b$ that marks a transition (with increasing $b$) from a $b$-range for which absorption is allowed to one for which it is forbidden. We call such a transition point (on the $b$ axis) a globally disconnecting point. In principle, there may also exist a point of transition in the other direction (namely, from a forbidden range to an allowed one upon increasing $b$) which we call a globally connecting point.